what are three consecutive odd numbers that add up to 357?

Answer by nyc_function(2741)

Let x = first odd number
Let x + 2 = second odd number
Let x + 4 = third odd number
Here is your equation:
x + x + 2 + x + 4 = 357
3x + 6 = 357
3x = 357 - 6
3x = 351
x = 351/3
x = 117
Here are the three consecutive odd numbers:
x = 117
x + 2 = 117 + 2 = 119
x + 4 = 117 + 4 = 121
They are: 117, 119, 121